# Global regularity for axisymmetric, swirl-free solutions of the Euler equation in four dimensions

### v1

@@ Abstract
In this paper, we prove global regularity for all smooth, axisymmetric, swirl-free solutions of the Euler equation in four dimensions. Previous works establishing global regularity for certain axisymmetric, swirl-free solutions of the Euler equation in four dimensions required the additional assumption that $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$, which can fail even for Schwartz class solutions. The key advance is a new bound on the vortex stretching term that only requires $\frac{\omega^{0}}{r^{2}}\in L^{2,1}(\mathbb{R}^{4})$, which is generically true for any axisymmetric, swirl-free initial data $u^{0}\in H^{s}\left(\mathbb{R}^{4}\right),s>4$, with reasonable decay at infinity.
@@ 1 Introduction
It is the difficulty of controlling the vortex stretching term that makes global regularity such a challenging open problem in three dimensions.
@@
This is related to the antisymmetric part of $\nabla u$ by
@@
Suppose $u^{0}\in H^{s}_{df}\left(\mathbb{R}^{4}\right),s>4$ is axisymmetric, swirl-free and that $\frac{\omega^{0}}{r^{2}}\in L^{2,1}\left(\mathbb{R}^{4}\right)$. Then there exists a global smooth solution of the Euler equation $u\in C\left([0,+\infty);H^{s}\left(\mathbb{R}^{4}\right)\right)$, and for all $0<t<+\infty$,
@@
where $C>0$ is an absolute constant independent of $u^{0}$. Furthermore, the condition $\frac{\omega^{0}}{r^{2}}\in L^{2,1}$ holds for all axisymmetric, swirl-free $u^{0}\in H^{s}_{df}\left(\mathbb{R}^{4}\right),s>4$ such that for all $r\geq 0,z\in\mathbb{R}$,
$|\omega^{0}(r,z)|\leq\frac{C}{(1+r^{2}+z^{2})^{2}}.$ (1.20)
@@
The regularity and decay assumptions are somewhat stronger than absolutely necessary, and could be relaxed somewhat. In particular, it the rate of decay at infinity could be optimized more, and the bound $\frac{\omega}{r}\in L^{\infty}$ which comes from the Sobolev embedding $H^{2+\epsilon}\hookrightarrow L^{\infty}$ could probably be relaxed as well. However, Theorem 1.1 completely resolves the global regularity problem for axisymmetric, swirl-free solutions of the Euler equation in four dimensions with smooth enough data and moderate decay at infinity.
All previous global regularity results for axisymmetric, swirl-free solutions of the Euler equation in four dimensions did not apply to generic smooth with decay at infinity, because they required that $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$, which can fail even for Schwarz class initial data. In particular Lim and Jeong proved global regularity for axisymmetric, swirl-free solutions of the Euler equation when $d=4$ if $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$ and $\omega^{0}$ is compactly supported [LimJeongARMA], giving the bound
@@
where $C_{\omega^{0}}>0$ is a constant depending on the initial data. This improved an earlier global regularity result—also assuming that $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$—proven by Choi, Jeong, and Lim [ChoiJeongLimProc], but with a bound that is exponential in time, rather than quadratic. This weaker result was also derived by the author and Tsai in [MillerTsai]. Furthermore, Lim proved global regularity in all dimensions $d\geq 3$ when $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$ under the additional assumption that $\omega^{0}$ has a single sign [LimSingleSign]. Note that both of these properties are preserved by the dynamics of the Euler equation.
We emphasize that the improvement from the requirement that $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$ in these previous papers to the requirement $\frac{\omega^{0}}{r^{2}}\in L^{2,1}$ is not merely technical, because the later holds for generic data in $H^{s}_{df}\left(\mathbb{R}^{4}\right),s>4$ with reasonable decay at infinity, while the former can fail for even Schwartz class initial data. For example, let
@@
The methods in this paper are somewhat related to the methods used by Danchin in [Danchin] to prove global regularity for axisymmetric swirl-free solutions of the Euler equation when $d=3$. Danchin proved global regularity when $d=3$ under the assumptions $\frac{\omega}{r}\in L^{3,1}$ and $\omega\in L^{\infty}\cap L^{3,1}$, which requires less regularity than the initial global regularity proof due to Yudovich, and which like the result in this paper has a key bound that involves controlling the vortex stretching term.
@@ 1.1 Definitions
We define the Sobolev space $H^{s}\left(\mathbb{R}^{4}\right),s\geq 0$ in terms of the Fourier transform.
@@
$\|f\|_{L^{p,\infty}}=\sup_{\tau>0}\left(\tau\mu\left(\left\{x:|f(x)|>\tau\right\}\right)^{\frac{1}{p}}\right)$ (1.28)
@@
The local wellposedness result was proven by Kato in [KatoLocalWP] and the Beale-Kato-Majda regularity criterion was proven in [BKM]. Note that it is common to denote the anti-symmetric part of $\nabla u$ by $\omega$. We use $A$ instead, because in our context this makes it clearer when we express the relationship between $A$ and the scalar vorticity $\omega(r,z)=\partial_{r}u_{z}-\partial_{z}u_{r}$.
@@ 3 Global regularity
In this section, we will show that the quasi-norm $\|\frac{\omega}{r^{2}}\|_{L^{2,1}}$ is a conserved quantity for the Euler equation. When combined with the bound on the rate of vortex stretching given in Theorem 2.5, this immediately leads to global regularity by providing an exponential bound on the growth of vorticity.
@@
Suppose $u\in C\left([0,T_{max});H^{s}\left(\mathbb{R}^{4}\right)\right),s>4$ is an axisymmetric, swirl-free solution of Euler’s equation, and that $\frac{\omega^{0}}{r^{2}}\in L^{2,1}\left(\mathbb{R}^{4}\right)$. Then for all $0<t<T_{max}$,
@@
$\omega\circ X=\omega^{0}.$ (3.4)
@@
Suppose $u^{0}\in H^{s}_{df}\left(\mathbb{R}^{4}\right),s>4$ is axisymmetric, swirl-free and $\frac{\omega^{0}}{r^{2}}\in L^{2,1}\left(r^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z\right)$. Then there exists a global smooth solution of the Euler equation $u\in C\left([0,+\infty);H^{s}\left(\mathbb{R}^{4}\right)\right)$, and for all $0<t<+\infty$,
@@ 4 Hypothesis satisfied for generic data
$u\in H^{s},s>4$ implies that $u\in C^{2}\left(\mathbb{R}^{4}\right)$, and consequently that $\frac{\omega}{r}\in L^{\infty}.$ See Proposition 3.22 in [MillerTsai] for details. This implies that for all $r\geq 0,z\in\mathbb{R}$
@@
and that consquently we have the bounds

### v2

@@ Abstract
In this paper, we prove global regularity for all smooth, axisymmetric, swirl-free solutions of the Euler equation in four dimensions. Previous works establishing global regularity for certain axisymmetric, swirl-free solutions of the Euler equation in four dimensions required the additional assumption that $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$, which can fail even for Schwartz class initial data. The key advance is a new bound on the vortex stretching term that only requires $\frac{\omega^{0}}{r^{2}}\in L^{2,1}(\mathbb{R}^{4})$, which is generically true for any axisymmetric, swirl-free initial data $u^{0}\in H^{s}\left(\mathbb{R}^{4}\right),s>4$, with reasonable decay at infinity.
@@ 1 Introduction
This equation can also be expressed equivalently as
$\partial_{t}\omega+(u\cdot\nabla)\omega-S\omega=0,$ (1.6)
where $S$ is the symmetric part of the velocity gradient, $S_{ij}=\frac{1}{2}(\partial_{i}u_{j}+\partial_{j}u_{i})$. It is the difficulty of controlling the vortex stretching term $S\omega$ that makes global regularity such a challenging open problem in three dimensions. Vortex stretching is also a feature in four and higher dimensions, and indeed is stronger relative to advection than in the three dimensional case. In this case the evolution of the vorticity is given by
$\partial_{t}A+(u\cdot\nabla)A+SA+AS=0.$ (1.7)
@@
This is related to the antisymmetric part of the velocity gradient $A$ by
@@
Suppose $u^{0}\in H^{s}_{df}\left(\mathbb{R}^{4}\right),s>4$ is axisymmetric, swirl-free and that $\frac{\omega^{0}}{r^{2}}\in L^{2,1}\left(\mathbb{R}^{4}\right)$. Then there exists a unique, global smooth solution of the Euler equation $u\in C\left([0,+\infty);H^{s}_{df}\left(\mathbb{R}^{4}\right)\right)$, and for all $0<t<+\infty$, this solution is axisymmetric, swirl-free and satisfies
@@
where $C>0$ is an absolute constant independent of $u^{0}$. Furthermore, the condition $\frac{\omega^{0}}{r^{2}}\in L^{2,1}$ holds for all axisymmetric, swirl-free $u^{0}\in H^{s}_{df}\left(\mathbb{R}^{4}\right),s>4$, such that for all $r\geq 0,z\in\mathbb{R}$,
$|\omega^{0}(r,z)|\leq\frac{C^{\prime}}{(1+r^{2}+z^{2})^{2}}.$ (1.22)
@@
The regularity and decay assumptions on the initial data in Theorem 1.1 are somewhat stronger than absolutely necessary and could be relaxed somewhat. In particular, the rate of decay at infinity could be optimized more, and the bound $\frac{\omega}{r}\in L^{\infty}$—which comes from the Sobolev embedding $H^{2+\epsilon}\hookrightarrow L^{\infty}$—could probably be relaxed as well. However, Theorem 1.1 completely resolves the global regularity problem for axisymmetric, swirl-free solutions of the Euler equation in four dimensions with smooth enough data and moderate decay at infinity.
All previous global regularity results for axisymmetric, swirl-free solutions of the Euler equation in four dimensions did not apply to generic smooth data with decay at infinity, because they required that $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$, which can fail even for Schwarz class initial data. In particular, Lim and Jeong proved global regularity for axisymmetric, swirl-free solutions of the Euler equation when $d=4$ if $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$ and $\omega^{0}$ is compactly supported [LimJeongARMA], giving the bound
@@
where $C_{\omega^{0}}>0$ is a constant depending on the initial data. This improved an earlier global regularity result—also assuming that $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$—proven by Choi, Jeong, and Lim [ChoiJeongLimProc], but with a bound that is exponential in time, rather than quadratic. This weaker result was also derived by the author and Tsai in [MillerTsai]. Furthermore, Lim proved global regularity in all dimensions $d\geq 3$ when $\frac{\omega^{0}}{r^{d-2}}\in L^{\infty}$ under the additional assumption that $\omega^{0}$ has a single sign [LimSingleSign]. Note that both of these properties are preserved by the dynamics of the Euler equation.
We emphasize that the improvement from the requirement that $\frac{\omega^{0}}{r^{2}}\in L^{\infty}$ in these previous papers, to the requirement $\frac{\omega^{0}}{r^{2}}\in L^{2,1}$ is not merely technical, because the later holds for generic data in $H^{s}_{df}\left(\mathbb{R}^{4}\right),s>4$ with reasonable decay at infinity, while the former can fail for even Schwartz class initial data. For example, let
@@
There are some similarities between the methods used in this paper to those used by Danchin in [Danchin] to prove global regularity for axisymmetric swirl-free solutions of the Euler equation when $d=3$. Danchin proved global regularity when $d=3$ under the assumptions $\frac{\omega}{r}\in L^{3,1}$ and $\omega\in L^{\infty}\cap L^{3,1}$, which requires less regularity than the initial global regularity proof due to Yudovich, and which, like the result in this paper, has a key bound that involves controlling the vortex stretching term.
The key difference in approach, beyond the fact that [Danchin] deals with $d=3$ rather than $d=4$, is that the current paper exploits the fact that the Biot-Savart law for axisymmetric, swirl-free solutions has a locally two dimensional structure, even in higher dimension. In particular in two dimensions, we have the Sobolev-type inequality:
$\|f\|_{L^{\infty}\left(\mathbb{R}^{2}\right)}\leq C\|\nabla f\|_{L^{2,1}\left(\mathbb{R}^{2}\right)}.$ (1.27)
Because $\frac{u_{r}}{r}$ and $\frac{\omega}{r}$ are both continuous functions and $\omega$ is a derivative of $u$, we might expect a bound of the form
$\left\|\frac{u_{r}}{r}\right\|_{L^{\infty}}\leq C\left\|\frac{\omega}{r}\right\|_{L^{2,1}(\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z)}.$ (1.28)
It is straightforward to show that this bound in fact holds from Proposition 2.3, but this ends up being less useful, because it is the physically meaningful measure $r^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z$, corresponding to the volume in $\mathbb{R}^{4}$, that is important. The bound that we will actually use is
$\left\|\frac{u_{r}}{r}\right\|_{L^{\infty}}\leq C\left\|\frac{\omega}{r^{2}}\right\|_{L^{2,1}\left(r^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z\right)}.$ (1.29)
This is roughly equivalent, because while the right hand sides of the bounds in (1.28) and (1.29) are not equivalent, if we replace the $L^{2,1}$ norm with the $L^{2}$ norm than these bounds are exactly equivalent with
$\left\|\frac{\omega}{r^{2}}\right\|_{L^{2}(r^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z)}^{2}=\int_{-\infty}^{\infty}\int_{0}^{\infty}\left|\frac{\omega(r,z)}{r^{2}}\right|^{2}r^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z=\int_{-\infty}^{\infty}\int_{0}^{\infty}\left|\frac{\omega(r,z)}{r}\right|^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z=\left\|\frac{\omega}{r}\right\|_{L^{2}(\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z)}^{2}.$ (1.30)
The space $L^{2,1}$ is just a slightly stronger version of $L^{2}$, so it is not surprising then that both (1.28) and (1.29) hold.
@@ 1.1 Definitions
We must define the Sobolev space $H^{s}$, as well as the Lorentz space $L^{p,q}$.
@@
$\|f\|_{L^{p,\infty}}=\sup_{\tau>0}\left(\tau\mu\left(\left\{x:|f(x)|>\tau\right\}\right)^{\frac{1}{p}}\right).$ (1.34)
Note that for all $1\leq p<+\infty$, we have
$\|f\|_{L^{p,p}}=\|f\|_{L^{p}}$ (1.35)
@@
The local wellposedness result was proven by Kato in [KatoLocalWP] and the Beale-Kato-Majda regularity criterion (1.41) was proven by Beale, Kato, and Majda for $d=3$ in [BKM] and by Kato and Ponce for $d\geq 4$ in [KatoPonce]. Note that it is common to denote the anti-symmetric part of $\nabla u$ by $\omega$ when $d\geq 4$, and the vorticity cannot be expressed by a vector. We use $A$ instead, because in our context this makes it clearer when we express the relationship between $A$ and the scalar vorticity $\omega(r,z)=\partial_{r}u_{z}-\partial_{z}u_{r}$.
@@ 3 Global regularity
In this section, we will show that the Lorentz quasi-norm $\|\frac{\omega}{r^{2}}\|_{L^{2,1}}$ is a conserved quantity for the Euler equation. When combined with the bound on the rate of vortex stretching given in Theorem 2.5, this immediately leads to global regularity by providing an exponential bound on the growth of vorticity in $L^{\infty}$.
@@
Suppose $u\in C\left([0,T_{max});H^{s}_{df}\left(\mathbb{R}^{4}\right)\right),s>4$ is an axisymmetric, swirl-free solution of Euler’s equation, and that $\frac{\omega^{0}}{r^{2}}\in L^{2,1}\left(\mathbb{R}^{4}\right)$. Then for all $0<t<T_{max}$,
@@
$\frac{\omega}{r^{2}}\circ X=\frac{\omega^{0}}{r^{2}}.$ (3.4)
@@
Suppose $u^{0}\in H^{s}_{df}\left(\mathbb{R}^{4}\right),s>4$ is axisymmetric, swirl-free and $\frac{\omega^{0}}{r^{2}}\in L^{2,1}\left(r^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z\right)$. Then there exists a unique, global smooth solution of the Euler equation $u\in C\left([0,+\infty);H^{s}_{df}\left(\mathbb{R}^{4}\right)\right)$, and for all $0<t<+\infty$, this solution is axisymmetric, swirl-free and satisfies
@@ 4 Hypothesis satisfied for generic data
The Sobolev embedding implies that $u\in C^{2}\left(\mathbb{R}^{4}\right)$, that $\nabla A\in C\left(\mathbb{R}^{4}\right)$ and consequently that $\frac{\omega}{r}\in L^{\infty}.$ See Proposition 3.22 in [MillerTsai] for details. This implies that for all $r\geq 0,z\in\mathbb{R}$
@@
and that consequently we have the bounds
@@
Remark 4.2.
Our result requires that $\frac{\omega^{0}}{r^{2}}\in L^{2,1}$, which is stronger than control in $L^{2}$. The key idea is that for the $L^{2}$ norm, which is just slightly weaker than the $L^{2,1}$ quasi-norm, we have
$\left\|\frac{\omega^{0}}{r^{2}}\right\|_{L^{2}\left(\mathbb{R}^{4}\right)}^{2}=4\pi\int_{-\infty}^{\infty}\int_{0}^{\infty}\left|\frac{\omega^{0}}{r^{2}}\right|^{2}r^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z=4\pi\int_{-\infty}^{\infty}\int_{0}^{\infty}\left|\frac{\omega^{0}}{r}\right|^{2}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}z.$ (4.13)
We know that $\frac{\omega^{0}}{r}$ must be bounded for smooth enough data, so it is straightforward to see that the integral in (4.13) is finite. It is then not surprising that the $L^{2,1}$ norm is also finite.